OpenStudy (anonymous):

The function f is continuous on the interval [3, 11] with some of its values given in the table below. Use the data in the table to approximate f ′(9).

OpenStudy (anonymous):

https://gyazo.com/34449492be7bc0f0156af969966447b3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm thinking they want you to use the mean value theorem to estimate the value of f ' (9)

OpenStudy (anonymous):

Never heard of that in my life.

OpenStudy (anonymous):

And this is my second time taking Calc

jimthompson5910 (jim_thompson5910):

one second while I think

OpenStudy (anonymous):

No problem boss

jimthompson5910 (jim_thompson5910):

ok I'm going to try something with the mean value theorem

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

\[\Large f \ ' (9) \approx \frac{f(10) - f(8)}{10 - 8}\] \[\Large f \ ' (9) \approx \frac{16 - 10}{10 - 8}\] \[\Large f \ ' (9) \approx \frac{6}{2}\] \[\Large f \ ' (9) \approx 3\] but none of the answer choices match that. But I guess it could easily be D

OpenStudy (anonymous):

I went ahead and put D. Thanks